# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

### v1

@@ 2.5 An estimate on the forcing term
    $\mathbb{R}\ni x\mapsto\frac{x^{2}}{x+e^{-x}-1}=:\Psi(x)\in\mathbb{R}$
    belongs to $C^{\infty}(\left[0,\infty\right[)$. Thus it is all okay for $\theta=2$ case, for the map $\xi\mapsto|\xi|^{2}$ is smooth in $\mathbb{R}_{\xi}^{N}$ and $C_{T}(\xi)=T^{-2}\Psi(T|\xi|^{2})$ , but the rest cases require additional care (see Appendix A).
@@
    Taking into account (2.5), we cannot obtain homogeneous Besov-type space version of the inequality (see also (A)). This is consistent with Theorem 1.5 (1) (we note that $B^{-\theta}_{(p,\infty),1}\setminus\dot{B}^{-\theta}_{(p,\infty),\infty}\neq\varnothing$).
@@
    The coefficient $C^{\flat}_{T}+C^{\sharp}_{T}$ appering in the proof of Proposition 2.16 satisfies
    $0<C^{\flat}_{T}+C^{\sharp}_{T}\leqslant CT^{-1}\quad\text{as }T\to 0.$
@@ 3 Proofs of main theorems
In this section, we show the existence of solution to (1.2) as a fixed point of the map defined below in Lorentz spaces. We carry out partially the same arguments as in [11], but we give a full proof for reader’s convenience.
To begin with, we fix $T>0$ and $p\in\left]1,\infty\right[$ and we write
$X_{T}:=\left\{u(x,t):\text{Lebesgue measurable in }\mathbb{R}^{N}\times\left]0,T\right[\mathrel{}\middle|\mathrel{}\left\|u\mid X_{T}\right\|<\infty\right\}$
, where
$\left\|u\mid X_{T}\right\|:=\sup_{0<t<T}\left\|u(\cdot,t)\mid L^{p,\infty}_{\mathrm{ul}}\right\|$.
@@
We take a constant $M>0$ (which is determined later by (3.2)) and set
@@ 3.1 Estimates on $I[\mu]$
In this subsection, we show two estimates on the operator $I[\mu]$ which correspond to Theorem 1.5 (2) and 1.5 (1). These kind of critical estimates were early obtained independently by Meyer [20] and Yamazaki [26] in Lorentz spaces. Recently, in homogeneous Besov spaces $\dot{B}^{s}_{p,\infty}$ [6] or in uniformly local weak Zygmund spaces [11], such critical estimate was obtained and played a key role.
@@
Fix $0<T<1$, $1<p<\infty$ and $r\in\left]1,p\right[$. Then $I[\mu]$ is a map from $B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}$ to $X_{T}$, and its operator norm is finite.
@@
We define $k(r)>0$ as $p^{-1}=(2r)^{-1}+(2k(r))^{-1}$. It is easy to see that $1<r<p<k(r)$. Then we let $\lambda>0$, set
$\eta=\eta(\lambda):=\lambda^{a},\quad a:=\frac{\theta pr}{2N(p-r)}$ (3.1)
@@
$I[\mu](x,t)$ $=\int_{0}^{t}S(\tau)\mu(x)\,d\tau=\int_{0}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\mu(x)\,d\tau$
$=\left(\int_{0}^{\eta}+\int_{\eta}^{\infty}\right)\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\mu(x)\,d\tau=:I_{\eta}^{0}[\mu](x,t)+I_{\eta}^{1}[\mu](x,t).$
@@
$\left\|I_{\eta}^{0}[\mu](x,t)\mid L^{r,\infty}_{\mathrm{ul}}\right\|\leqslant C\int_{0}^{\eta}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\mu(x)\mathrel{}\middle|\mathrel{}B^{0}_{(r,\infty),1}\right\|\,d\tau$
$\leqslant C\int_{0}^{\eta}\tau^{-1+\frac{N}{\theta r}-\frac{N}{\theta p}}\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|=C\lambda^{1/2}\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|$
@@
$\left\|I_{\eta}^{1}[\mu](x,t)\mathrel{}\middle|\mathrel{}L^{k(r),\infty}_{\mathrm{ul}}\right\|\leqslant C\int_{\eta}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\mu(x)\mathrel{}\middle|\mathrel{}B^{0}_{(k(r),\infty),1}\right\|\,d\tau$
$\leqslant C\int_{\eta}^{\infty}\tau^{-1+\frac{N}{\theta k(r)}-\frac{N}{\theta p}}\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{k(r)}-\frac{N}{p}}_{(k(r),\infty),\infty}\right\|=C\lambda^{-1/2}\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{k(r)}-\frac{N}{p}}_{(k(r),\infty),\infty}\right\|.$
Now we use Corollary 2.8 in the form $L^{p,\infty}_{\mathrm{ul}}\supset\left(L^{r,\infty}_{\mathrm{ul}},L^{k(r),\infty}_{\mathrm{ul}}\right)_{1/2,\infty}$ to derive
$\left\|I[\mu](x,t)\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\sup_{\lambda>0}\lambda^{-1/2}\inf_{I[\mu]=I_{0}+I_{1}}\left\{\left\|I_{0}\mathrel{}\middle|\mathrel{}L^{r,\infty}_{\mathrm{ul}}\right\|+\lambda\left\|I_{1}\mathrel{}\middle|\mathrel{}L^{k(r),\infty}_{\mathrm{ul}}\right\|\right\}$
$\leqslant C\sup_{\lambda>0}\left\{\lambda^{-1/2}\left\|I_{\eta(\lambda)}^{0}[\mu]\mathrel{}\middle|\mathrel{}L^{r,\infty}_{\mathrm{ul}}\right\|+\lambda^{1/2}\left\|I_{\eta(\lambda)}^{1}[\mu]\mathrel{}\middle|\mathrel{}L^{k(r),\infty}_{\mathrm{ul}}\right\|\right\}$
$\leqslant C\left\{\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|+\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{k(r)}-\frac{N}{p}}_{(k(r),\infty),\infty}\right\|\right\}$
$\leqslant C\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|$
for all $t<T$. In the last inequality, we used Proposition 2.10. ∎
@@
Fix $0<T<1$. Let $1<p<\infty$. Then $I[\mu]$ is a map from $B^{-\theta}_{(p,\infty),1}$ to $X_{T}$, and its operator norm is finite.
@@
We set
$\Psi_{j}(x):=\mathcal{F}^{-1}\left[\Phi_{j}\right](x)\,\,\text{for }j\in\mathbb{Z},\quad\Psi_{(0)}(x):=\mathcal{F}^{-1}\left[\Phi_{(0)}\right](x)$
, where $\Phi_{j}$ and $\Phi_{(0)}$ are the same as in the proof of Proposition 2.16. By using Proposition 2.3 and 2.9, we have
$\left\|\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant C\left\|\int_{0}^{t}S(\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{(p,\infty),1}\right\|$
$=C\sum_{j\in\mathbb{N}}2^{0j}\left\|\Delta_{j}\int_{0}^{t}S(\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$=C\sum_{j\in\mathbb{N}}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\ast\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant C\sum_{j\in\mathbb{N}}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$. (3.2)
Fix $j\in\mathbb{N}$ and sufficiently small $r>0$. Because $\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}L^{1}\right\|$ is independent of $j\in\mathbb{N}$, Proposition 2.12 and 2.13 give that
$\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|\leqslant\sum_{l=-2}^{2}2^{\theta(j+l)}2^{-\theta(j+l)}\left\|\Delta_{j+l}\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|$
@@
$\leqslant C2^{-\theta j}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}B^{\theta}_{1,\infty}\right\|$
$=C2^{-\theta j}\left\|\int_{0}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}\left(B^{\theta-\theta r}_{1,\infty},B^{\theta+\theta r}_{1,\infty}\right)_{1/2,\infty}\right\|$
$\leqslant C2^{-\theta j}\sup_{\lambda>0}\lambda^{-1/2}\bigg\{\int_{0}^{\lambda^{1/2r}}\chi_{\left\{0<\tau<t\right\}}(\tau)\left(1+\tau^{r-1}\right)\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}B^{0}_{1,\infty}\right\|\,d\tau$
$\qquad\qquad\qquad\qquad\quad+\lambda\int_{\lambda^{1/2r}}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)\left(1+\tau^{-1-r}\right)\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}B^{0}_{1,\infty}\right\|\,d\tau\bigg\}$
$\leqslant C2^{-\theta j}\sup_{\lambda>0}\lambda^{-1/2}\bigg\{\int_{0}^{\lambda^{1/2r}}\tau^{r-1}\,d\tau+\lambda\int_{\lambda^{1/2r}}^{\infty}\tau^{-1-r}\,d\tau\bigg\}\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}L^{1}\right\|=C2^{-\theta j}$
holds for $0<t\leqslant T<1$ (Here we wrote $\Delta_{j}:=\Delta_{0}$ for $j<0$). Plugging this estimate in (3.1) and taking the supremum over $t\in\left]0,T\right[$ yield the claim. ∎

### v2

@@ 2.5 An estimate on the inhomogeneous term
    $\mathbb{R}\ni x\mapsto\Psi(x):=\frac{x^{2}}{x+e^{-x}-1}\in\mathbb{R}$
    belongs to $C^{\infty}(\left[0,\infty\right[)$. Thus $\theta\in 2\mathbb{Z}_{\geqslant 1}$ case are okay, for the map $\xi\mapsto|\xi|^{2k}$ is smooth in $\mathbb{R}_{\xi}^{N}$ for all $k\in\mathbb{Z}_{\geqslant 1}$ and $C_{T}(\xi)=T^{-2}\Psi(T|\xi|^{2})$ , but the rest cases require additional care (see Appendix A).
@@
    Taking into account (2.5), we cannot obtain homogeneous Besov-type space version of the inequality (see also Remark A.1). This is consistent with Theorem 1.5 (1) (we note that $B^{-\theta,1}_{p,q,\infty}\setminus\dot{B}^{-\theta,\infty}_{p,q,\infty}\neq\varnothing$).
@@
    The coefficient $C^{\flat}_{T}+C^{\sharp}_{T}$ appearing in the proof of Proposition 2.18 satisfies
    $(0<)\,C^{\flat}_{T}+C^{\sharp}_{T}\leqslant CT^{-1}\quad\text{for }T\sim 0.$
@@ 3 Proofs of main theorems
In this section, we show the existence of solution to (1.1) as a fixed point of the map defined below. To begin with, we fix $T>0$ and $1<q\leqslant p<\infty$ and we write
$X_{T}:=\left\{u(x,t):\text{Lebesgue measurable in }\mathbb{R}^{N}\times\left]0,T\right[\mathrel{}\middle|\mathrel{}\left\|u\mid X_{T}\right\|<\infty\right\},$
where
$\left\|u\mid X_{T}\right\|:=\sup_{0<t<T}\left\|u(\cdot,t)\mid M^{p}_{q,\infty}\right\|.$
@@
We take a constant $M>0$ (which is determined later by (3.2) and (3.7)) and set
@@ 3.1 Estimates on $I[\mu]$
In this subsection, we show two estimates on the operator $I[\mu]$ which correspond to Theorem 1.5 (2) and 1.5 (1). These kind of critical estimates were early obtained independently by Meyer [24] and Yamazaki [29] in Lorentz spaces. Recently, in homogeneous Besov spaces $\dot{B}^{s}_{p,\infty}$ [6], in the predual of (nonlocal) Lorentz–Morrey spaces $\mathcal{M}^{p}_{q,\infty}$ [8] or in uniformly local weak Zygmund spaces [14], such critical estimate was obtained and played a key role.
@@
Fix $0<T<1$, $1<q\leqslant p<\infty$ and
$r\in\left]\max\left\{\frac{p}{q};\frac{Np}{N+\theta p}\right\},p\right[.$
Then $I[\mu]$ is a map from $B^{N/r-N/p-\theta,\infty}_{r,rq/p,\infty}$ to $X_{T}$, and its operator norm is finite.
@@
We take $p_{0}\in\left[r,p\right[$ so that $q_{0}:=p_{0}q/p>1$ and $2p_{0}>p$ hold. Then we put $p_{1}:=pp_{0}/(2p_{0}-p)$, $q_{1}:=qp_{0}/(2p_{0}-p)$ ($\Rightarrow p_{1}\geqslant q_{1}>1$). We let $\lambda>0$, set
$\eta=\eta(\lambda):=\lambda^{a},\quad a:=\frac{\theta pp_{0}}{2N(p-p_{0})}$ (3.1)
@@
$I[\mu](x,t)$ $=\int_{0}^{t}S(\tau)\mu(x)\,d\tau=\int_{0}^{\infty}\chi_{\left]0,t\right[}(\tau)S(\tau)\mu(x)\,d\tau$
$=\left(\int_{0}^{\eta}+\int_{\eta}^{\infty}\right)\chi_{\left]0,t\right[}(\tau)S(\tau)\mu(x)\,d\tau=:I_{\lambda}^{0}[\mu](x,t)+I_{\lambda}^{1}[\mu](x,t).$
@@
$\left\|I_{\lambda}^{0}[\mu](x,t)\mathrel{}\middle|\mathrel{}M^{p_{0}}_{q_{0},\infty}\right\|\leqslant C\int_{0}^{\eta}\chi_{\left]0,t\right[}(\tau)\left\|S(\tau)\mu(x)\mathrel{}\middle|\mathrel{}B^{0,1}_{p_{0},q_{0},\infty}\right\|\,d\tau$
$\leqslant C\int_{0}^{\eta}\tau^{-1+\frac{N}{\theta p_{0}}-\frac{N}{\theta p}}\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{p_{0}}-\frac{N}{p}-\theta,\infty}_{p_{0},q_{0},\infty}\right\|=C\lambda^{1/2}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{p_{0}}-\frac{N}{p}-\theta,\infty}_{p_{0},q_{0},\infty}\right\|$
@@
$\left\|I_{\lambda}^{1}[\mu](x,t)\mathrel{}\middle|\mathrel{}M^{p_{1}}_{q_{1},\infty}\right\|\leqslant C\int_{\eta}^{\infty}\chi_{\left]0,t\right[}(\tau)\left\|S(\tau)\mu(x)\mathrel{}\middle|\mathrel{}B^{0,1}_{p_{1},q_{1},\infty}\right\|\,d\tau$
$\leqslant C\int_{\eta}^{\infty}\tau^{-1+\frac{N}{\theta p_{1}}-\frac{N}{\theta p}}\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{p_{1}}-\frac{N}{p}-\theta}_{p_{1},q_{1},\infty}\right\|=C\lambda^{-1/2}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{p_{1}}-\frac{N}{p}-\theta}_{p_{1},q_{1},\infty}\right\|.$
Now we use Proposition 2.8 in the form $M^{p}_{q,\infty}\supset\left(M^{p_{0}}_{q_{0},\infty},M^{p_{1}}_{q_{1},\infty}\right)_{1/2,\infty}$ to derive
$\left\|I[\mu](x,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $\leqslant C\sup_{\lambda>0}\lambda^{-1/2}\inf_{I[\mu]=I_{0}+I_{1}}\left\{\left\|I_{0}\mathrel{}\middle|\mathrel{}M^{p_{0}}_{q_{0},\infty}\right\|+\lambda\left\|I_{1}\mathrel{}\middle|\mathrel{}M^{p_{1}}_{q_{1},\infty}\right\|\right\}$
$\leqslant C\sup_{\lambda>0}\left\{\lambda^{-1/2}\left\|I_{\lambda}^{0}[\mu]\mathrel{}\middle|\mathrel{}M^{p_{0}}_{q_{0},\infty}\right\|+\lambda^{1/2}\left\|I_{\lambda}^{1}[\mu]\mathrel{}\middle|\mathrel{}M^{p_{1}}_{q_{1},\infty}\right\|\right\}$
$\leqslant C\left\{\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{p_{0}}-\frac{N}{p}-\theta,\infty}_{p_{0},q_{0},\infty}\right\|+\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{p_{1}}-\frac{N}{p}-\theta,\infty}_{p_{1},q_{1},\infty}\right\|\right\}$
$\leqslant C\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{p_{0}}-\frac{N}{p}-\theta,\infty}_{p_{0},q_{0},\infty}\right\|\leqslant C\left\|\mu\mathrel{}\middle|\mathrel{}B^{\frac{N}{r}-\frac{N}{p}-\theta,\infty}_{r,\frac{rq}{p},\infty}\right\|$
for all $t<T<1$. In the last line, we used Proposition 2.12. ∎
@@
Fix $0<T<1$. Let $1<q\leqslant p<\infty$. Then $I[\mu]$ is a map from $B^{-\theta,1}_{p,q,\infty}$ to $X_{T}$, and its operator norm is finite.
@@
We set $\Psi_{j}(x):=\mathcal{F}^{-1}\left[\Phi_{j}\right](x)$ for $j\in\mathbb{Z}_{\geqslant 1}$ and $\Psi_{0}(x):=\mathcal{F}^{-1}\left[\Phi_{(0)}\right](x)$, where $\Phi_{j}$ and $\Phi_{(0)}$ are the same as in the proof of Lemma 2.13. Let $t\in\left]0,T\right[$. By using Proposition 2.3 and 2.11, we have
$\left\|\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant C\sum_{j\in\mathbb{N}}\left\|\Delta_{j}\int_{0}^{t}S(\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$=C\sum_{j\in\mathbb{N}}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\ast\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant C\sum_{j\in\mathbb{N}}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|.$ (3.2)
Because $\mathrm{supp}\,\Phi_{j}\subset\mathrm{supp}\left[\sum_{l=-2}^{2}\varphi_{j+l}\right]$ holds for $j\in\mathbb{Z}_{\geqslant 1}$, we have
$\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|\leqslant\sum_{l=-2}^{2}2^{-\theta(j+l)}2^{\theta(j+l)}\left\|\Delta_{j+l}\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|$
@@
$\leqslant C2^{-\theta j}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}B^{\theta,\infty}_{1}\right\|$
$\leqslant C2^{-\theta j}\left\|\int_{0}^{\infty}\chi_{\left]0,t\right[}(\tau)S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}\left(B^{\theta-\theta r,\infty}_{1},B^{\theta+\theta r,\infty}_{1}\right)_{1/2,\infty}\right\|.$
Here we wrote $\Delta_{j}:=\Delta_{0}$ and $\Psi_{j}:=\Psi_{0}$ for $j<0$, and we used Proposition 2.14 in the last for sufficiently small $r>0$. Because $\left\|\Psi_{j}\mathrel{}\middle|\mathrel{}L^{1}\right\|$ is independent of $j\in\mathbb{N}$, the decay estimate of $S(\tau)$ on the Besov spaces [1, Theorem 3] reveals that
$\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|$
$\leqslant C2^{-\theta j}\sup_{\lambda>0}\lambda^{-1/2}\bigg\{\int_{0}^{\lambda^{1/2r}}\chi_{\left]0,t\right[}(\tau)\left(1+\tau^{r-1}\right)\left\|\Psi_{j}\mathrel{}\middle|\mathrel{}B^{0,\infty}_{1}\right\|\,d\tau$
$\qquad\qquad\qquad\qquad\quad+\lambda\int_{\lambda^{1/2r}}^{\infty}\chi_{\left]0,t\right[}(\tau)\left(1+\tau^{-1-r}\right)\left\|\Psi_{j}\mathrel{}\middle|\mathrel{}B^{0,\infty}_{1}\right\|\,d\tau\bigg\}$
$\leqslant C2^{-\theta j}\sup_{\lambda>0}\lambda^{-1/2}\bigg\{\int_{0}^{\lambda^{1/2r}}\tau^{r-1}\,d\tau+\lambda\int_{\lambda^{1/2r}}^{\infty}\tau^{-1-r}\,d\tau\bigg\}\left\|\Psi_{j}\mathrel{}\middle|\mathrel{}L^{1}\right\|=C2^{-\theta j}$
holds for each $j\in\mathbb{N}$. Plugging this estimate in (3.1) and taking the supremum over $t\in\left]0,T\right[$ yield the claim. ∎